# Distributionally Robust Markov Games with Average Reward

### v1

@@ Appendix D Proof of Section 6
Such an assumption is necessary for convergence Li (2003); Hu and Wellman (2003); Littman et al. (2001); Bowling (2001, 2000).
Theorem D.2.
Additionally assume Assumption D.1. The robust Nash-Iteration algorithm converge to an average reward Nash Equilibrium of the distributionally robust game.
@@
It then remains to show the convergence of Algorithm 1, and it is sufficient to show that for any $i$, the operator $\mathbf{T}=(\mathbf{T}_{1},...,\mathbf{T}_{N}):\mathbb{R}^{NS}\to\mathbb{R}^{NS}$ is a contraction under Span semi-norm, where
@@
here $\pi$ is the Nash Equilibrium specified by Algorithm 1. Consider $h,h^{\prime}$, and we consider the two cases from Assumption D.1. If $\pi$ is global optimal, then
@@
And if the second case from Assumption D.1 holds, then
@@
Thus Lemma G.2 implies that
@@
for some irreducible transition matrix $\mathsf{P}$. Applying Lemma G.1 implies
@@
Now since for irreducible transition matrix $\mathsf{P}$, $\alpha(\mathsf{P})<1$ Puterman (1994), then
@@
Appendix E Proof of Section 7
Theorem E.1.
@@ Appendix E Proof of Section 7
(1). By the uniform convergence of the robust discounted value function derived in Wang et al. (2023a), we have that
@@
The proof is a direct corollary of Theorem 7.1 and results from Roch, Atia, and Wang (2025). Specifically, it is shown in Theorem 3.4 in Roch, Atia, and Wang (2025) that, for any $\gamma>1-\frac{\epsilon}{D}$ [^3] The results in Roch, Atia, and Wang (2025) is derived for $\mathcal{H}$ which is the robust optimal span. However, their results also hold for any constant that is larger than $\mathcal{H}$. Our claim thus holds by noting that $D\geq\mathcal{H}$ Roch, Atia, and Wang (2025)., it holds that
@@
Appendix F Extension to Unichain Cases
In this section, we mainly discuss the extension of the solvability results to the unichain cases. The remaining results can be derived following a similar way, by re-deriving the results based on the multi-step operator introduced below.
Definition F.1 (Robust Bellman Operators).
For a fixed policy $\pi$, the robust transition operator $\mathcal{T}^{\pi}_{\mathcal{P}}:\mathbb{R}^{|\mathcal{S}|}\to\mathbb{R}^{|\mathcal{S}|}$ is defined as:
$(\mathcal{T}_{\mathcal{P}_{s}^{\pi(s)}}V)(s):=\min_{\mathsf{P}\in\mathbb{P}_{s}^{\pi(s)}}\sum_{s^{\prime}\in\mathcal{S}}\mathsf{P}(s^{\prime}|s,\pi(s))V(s^{\prime})$
The robust Bellman evaluation operator $\hat{\mathcal{T}}^{\pi}_{\mathcal{P}}:\mathbb{R}^{|\mathcal{S}|}\to\mathbb{R}^{|\mathcal{S}|}$ is defined as:
$(\hat{\mathcal{T}}_{\mathcal{P}^{\pi(s)}_{s}}V)(s):=r(s,\pi(s))+(\mathcal{T}_{\mathcal{P}^{\pi(s)}_{s}}V)(s)$
where $r(s,\pi(s))$ is the reward and $\mathbb{P}^{a}_{s}$ is the uncertainty set of transition kernels for state-action pair $(s,a)$.
We first derive the following lemma by extending from the non-robust results from Puterman (1990).
Lemma F.2 (Multi-Step Span-Contraction).
Assume the robust unichain assumption holds for the policy $\pi$. Then the operator $\mathcal{T}_{\mathcal{P}}^{\pi}$ is a multi-step span-contraction under some additional technique assumptions. That is, there exists an integer $k\geq 1$ and a modulus $\gamma\in[0,1)$ such that for any vectors $V,U\in\mathbb{R}^{|\mathcal{S}|}$:
$\text{Span}((\mathcal{T}_{\mathcal{P}}^{\pi})^{k}(V)-(\mathcal{T}_{\mathcal{P}}^{\pi})^{k}(U))\leq\gamma\cdot\text{Span}(V-U)$
Proof.
The proof is derived similarly to the Theorem 8.5.2 of Puterman (1994). ∎
Corollary F.3.
The operator $\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}$ inherits the same multi-step span-contraction property, since for any $V,U$:
$\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V)-\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(U)=(r^{\pi}+\mathcal{T}_{\mathcal{P}}^{\pi}(V))-(r^{\pi}+\mathcal{T}_{\mathcal{P}}^{\pi}(U))=\mathcal{T}_{\mathcal{P}}^{\pi}(V)-\mathcal{T}_{\mathcal{P}}^{\pi}(U)$
Therefore, $\text{Span}((\hat{\mathcal{T}}_{\mathcal{P}}^{\pi})^{k}(V)-(\hat{\mathcal{T}}_{\mathcal{P}}^{\pi})^{k}(U))\leq\gamma\cdot\text{Span}(V-U)$.
We now state and prove the main theorem.
Theorem F.4.
Under the robust unichain assumption for a fixed policy $\pi$, the robust average-reward Bellman equation (5) has a solution.
Proof.
We establish existence by analyzing the convergence of the Relative Value Iteration (RVI) algorithm. Let $V_{0}\in\mathbb{R}^{|\mathcal{S}|}$ be an arbitrary initial vector, and define the sequence $\{V_{n}\}$ for $n\geq 0$ as:
$V_{n+1}=\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n})$
Let $u_{n}=V_{n+1}-V_{n}$. We first show that $\text{Span}(u_{n})\to 0$.
$\text{Span}(u_{n+1})=\text{Span}(V_{n+2}-V_{n+1})=\text{Span}(\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n+1})-\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n}))$
By applying Corollary F.3 repeatedly, we get:
$\text{Span}(u_{n+k})=\text{Span}(\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n+k-1})-\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n+k-2}))\leq\gamma\cdot\text{Span}(V_{n+k-1}-V_{n+k-2})=\gamma\cdot\text{Span}(u_{n+k-1})$
This seems incorrect. Let’s apply the k-step property correctly.
$\text{Span}(u_{n+k})=\text{Span}(\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n+k-1})-\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n})\text{ term})$
Let’s re-evaluate.
$\text{Span}(u_{n+k})=\text{Span}(V_{n+k+1}-V_{n+k})=\text{Span}(\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n+k})-\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n+k-1}))$
Let’s analyze $V_{n+k}-V_{n+k-1}$ as a function of earlier differences. $\text{Span}(V_{n+k}-V_{n+k-1})=\text{Span}(\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n+k-1})-\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n+k-2}))$. Let’s analyze $\text{Span}(V_{m+k}-V_{m})$. The key consequence of Lemma F.2 is that the sequence of differences $u_{n}$ approaches a constant vector. That is, $\lim_{n\to\infty}\text{Span}(u_{n})=0$. This means that for large $n$, $V_{n+1}(s)-V_{n}(s)\approx g$ for all $s\in\mathcal{S}$.
(a) Convergence of the Gain. The property $\lim_{n\to\infty}\text{Span}(V_{n+1}-V_{n})=0$ ensures that the sequence of scalars $\{g_{n}\}$, defined by picking an arbitrary reference state $s_{0}\in\mathcal{S}$,
$g_{n}:=V_{n+1}(s_{0})-V_{n}(s_{0})$
converges to a limit, which we denote by $g^{\pi}=\lim_{n\to\infty}g_{n}$.
(b) Convergence of the Differential Value Function. Define the differential value function $\tilde{V}_{n}=V_{n}-V_{n}(s_{0})\mathbf{1}$. This sequence lies in the subspace of vectors $V$ where $V(s_{0})=0$. It can be shown that the operator $B(V)=\hat{\mathcal{T}}^{\pi}_{\mathcal{P}}(V)-(\hat{\mathcal{T}}_{\mathcal{P}^{\pi(s_{0})}_{s_{0}}}(V))(s_{0})\mathbf{1}$ is a $k$-step contraction in the max-norm $||\cdot||_{\infty}$ on this subspace. The sequence $\tilde{V}_{n}$ can be shown to be a Cauchy sequence. Therefore, $\{\tilde{V}_{n}\}$ converges to a limit $V^{\pi}\in\mathbb{R}^{|\mathcal{S}|}$ with $V^{\pi}(s_{0})=0$.
(c) The Limit is a Solution. We have $V_{n}=\tilde{V}_{n}+V_{n}(s_{0})\mathbf{1}$. The RVI update is $V_{n+1}=\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n})$. Let’s analyze the update for $\tilde{V}_{n+1}$:
$\tilde{V}_{n+1}+V_{n+1}(s_{0})\mathbf{1}$ $=V_{n+1}=\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n})=\hat{\mathcal{T}}_{\mathcal{P}^{\pi(s_{0})}_{s_{0}}}(\tilde{V}_{n}+V_{n}(s_{0})\mathbf{1})$
$=\hat{\mathcal{T}}_{\mathcal{P}^{\pi(s_{0})}_{s_{0}}}^{\pi}(\tilde{V}_{n})+V_{n}(s_{0})\mathbf{1}\quad(\text{using shift invariance }\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V+c\mathbf{1})=\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V)+c\mathbf{1})$
Rearranging terms, we get:
$\tilde{V}_{n+1}+(V_{n+1}(s_{0})-V_{n}(s_{0}))\mathbf{1}=\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(\tilde{V}_{n})$
$\tilde{V}_{n+1}+g_{n}\mathbf{1}=\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(\tilde{V}_{n})$
. Taking the limit as $n\to\infty$, and using the continuity of the operator $\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}$:
$\lim_{n\to\infty}\tilde{V}_{n+1}+\lim_{n\to\infty}g_{n}\mathbf{1}=\lim_{n\to\infty}\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(\tilde{V}_{n})=\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(\lim_{n\to\infty}\tilde{V}_{n})$
$V^{\pi}+g^{\pi}\mathbf{1}=\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V^{\pi})$
This confirms the existence of a solution $(g^{\pi}_{\mathcal{P}},V^{\pi}_{\mathcal{P}})$.
Appendix G Technical Lemmas
Lemma G.1.
Define the ergodicity coefficient of a stochastic matrix $\mathsf{P}$ as
Proof.
For any two states $i,j$:
$=(1-\alpha(\mathsf{P}))\text{sp}(V)$
Lemma G.2 (Difference of Robust Operators).
For any $V_{1},V_{2}\in\mathbb{R}^{|\mathcal{S}|}$, there exists a valid transition kernel $\mathsf{P}^{*}$ constructible from $\mathcal{P}$ such that $\sigma^{\pi}(V_{1})-\sigma^{\pi}(V_{2})=\mathsf{P}^{*}(V_{1}-V_{2})$.
Proof.
Since $\mathcal{P}_{s}^{a}$ is a convex set and the function $h(\mathsf{P})=\mathsf{P}(V_{1}-V_{2})$ is continuous in $\mathsf{P}$, by the Intermediate Value Theorem there must exist a $\mathsf{P}_{s,a}^{*}\in\mathcal{P}_{s}^{a}$ such that $\sigma_{\mathcal{P}_{s}^{a}}(V_{1})-\sigma_{\mathcal{P}_{s}^{a}}(V_{2})=\mathsf{P}_{s,a}^{*}(V_{1}-V_{2})$. We then construct the global kernel $\mathsf{P}^{*}$ row by row as $\mathsf{P}^{*}(s^{\prime}|s)=\sum_{a}\pi(a|s)\mathsf{P}_{s,a}^{*}(s^{\prime}|s,a)$. This leads to the desired vector equality. ∎

### v2

@@ Appendix E Proof of Section 6
Such an assumption is necessary for convergence (li2003learning; hu2003nash; littman2001friend; Bowling2001; bowling2000convergence).
Theorem E.1.
Additionally assume Assumption 3. The robust Nash-Iteration algorithm converge to an average reward Nash Equilibrium of the distributionally robust game.
@@
It then remains to show the convergence of Algorithm 1, and it is sufficient to show that for any $i$, the operator $\mathbf{T}=(\mathbf{T}_{1},...,\mathbf{T}_{N}):\mathbb{R}^{NS}\to\mathbb{R}^{NS}$ is a contraction under Span heminorm, where
@@
here $\pi$ is the Nash Equilibrium specified by Algorithm 1. Consider $h,h^{\prime}$, and we consider the two cases from Assumption 3. If $\pi$ is global optimal, then
@@
And if the second case from Assumption 3 holds, then
@@
Thus Lemma C.7 implies that
@@
for some irreducible transition matrix $\mathsf{P}$. Applying Lemma C.6 implies
@@
Now since for irreducible transition matrix $\mathsf{P}$, $\alpha(\mathsf{P})<1$ (puterman1994markov), then
@@
Appendix F Proof of Section 7
Theorem F.1.
@@ Appendix F Proof of Section 7
(1). By the uniform convergence of the robust discounted value function derived in (wang2023robust), we have that
@@
The proof is a direct corollary of Theorem 7.1 and results from (Roch2025reduction). Specifically, it is shown in Theorem 3.4 in (Roch2025reduction) that, for any $\gamma>1-\frac{\epsilon}{D}$ [^2] The results in (Roch2025reduction) is derived for $\mathcal{H}$ which is the robust optimal span. However, their results also hold for any constant that is larger than $\mathcal{H}$. Our claim thus holds by noting that $D\geq\mathcal{H}$ (Roch2025reduction)., it holds that